OpenStudy (anonymous):

Consider the following expression: 4y+6y^2+7 ___________ 2+3y Divide the polynomial in the numerator by the binomial in the denomiator

myininaya (myininaya):

2y --------------- 3y+2| 6y^2+4y+7 -(6y^2+4y) ------------ 7

OpenStudy (anonymous):

and how is this to be written?

myininaya (myininaya):

\[\frac{6y^2+4y+7}{3y+2}=2y+\frac{7}{3y+2}\]